# GETS: Ensemble Temperature Scaling for Calibration in Graph Neural Networks

### v1

@@ Abstract
Graph Neural Networks (GNNs) deliver strong classification results but often suffer from poor calibration performance, leading to overconfidence or underconfidence. This is particularly problematic in high-stakes applications where accurate uncertainty estimates are essential. Existing post-hoc methods, such as temperature scaling, fail to effectively utilize graph structures, while current GNN calibration methods often overlook the potential of leveraging diverse input information and model ensembles jointly. In the paper, we propose Graph Ensemble Temperature Scaling (GETS), a novel calibration framework that combines input and model ensemble strategies within a Graph Mixture-of-Experts (MoE) architecture. GETS integrates diverse inputs, including logits, node features, and degree embeddings, and adaptively selects the most relevant experts for each node’s calibration procedure. Our method outperforms state-of-the-art calibration techniques, reducing expected calibration error (ECE) by $\geq 25\%$ across 10 GNN benchmark datasets. Additionally, GETS is computationally efficient, scalable, and capable of selecting effective input combinations for improved calibration performance. Codes are available when published.
00footnotetext: Equal contribution. 🖂Corresponding author.
@@ 1 Introduction
Graph Neural Networks (GNNs) have emerged as powerful tools for learning representations in numerous real-world applications, including social networks, recommendation systems, biological networks, and traffic systems, achieving state-of-the-art performance in tasks like node classification, link prediction, and graph classification (Kipf & Welling, 2016; Veličković et al., 2017). Their ability to capture complex relationships and dependencies makes them invaluable for modeling and predicting interconnected systems (Fan et al., 2019; Wu et al., 2019; Scarselli et al., 2008; Wu et al., 2022; 2021).
However, ensuring the reliability and trustworthiness of GNN prediction remains a critical challenge, especially in high-stakes applications where decisions can have significant consequences (e.g. human safety). Take GNN classification as an example, one key aspect of model reliability refinement is uncertainty calibration, which aims to align the predicted probabilities with the true likelihood of outcomes (Guo et al., 2017). In short, well-calibrated models provide confidence estimates that reflect real-world probabilities, which is essential for risk assessment and informed decision-making (Zhuang et al., 2023).
Calibration methods are typically applied post-hoc and fall into two categories: nonparametric and parametric. Nonparametric methods, like histogram binning and isotonic regression, adjust confidence estimates based on observed data distributions without assuming a specific functional form (Zadrozny & Elkan, 2001; Naeini et al., 2015). While allowing for the freedom in modeling, these methods often require large calibration datasets and struggle with high-dimensional or unevenly distributed graph data, making them less practical for graphs with vast numbers of nodes. In contrast, parametric methods such as temperature scaling (TS) and Platt scaling assume a functional relationship to adjust model outputs (Guo et al., 2017; Platt et al., 1999). Among them, TS has gained popularity for its simplicity and effectiveness in calibrating models by smoothing predicted probabilities. This method is scalable, and easy to implement, especially for large-scale graphs. Moving beyond the uniform adjustments applied to all data in classic TS, recent work focuses on learning node-specific calibration parameters to capture individual characteristics that impact graph calibration performance (Tang et al., 2024; Hsu et al., 2022).
@@
Given the challenges we have, we aim to resolve the following issues: (i) What are the influential factors in the calibration task for GNN classification? (ii) How can we wisely integrate these factors, which may exist in different spaces and scales, to effectively perform GNN calibration?
Figure 1: Expected calibration error (ECE), see Equation 3, for a well-trained CaGCN model (Wang et al., 2021). ECE is measured grouping nodes based on degree ranges rather than predicted confidence. Datasets are sorted by average degree $\frac{2|\mathcal{E}|}{|\mathcal{V}|}$ as a measure of connectivity (see Table 1), from low to high. In lower-connectivity datasets like Pubmed and Cora-full, high-degree nodes tend to show larger calibration errors, whereas in more connected datasets like Photo and Ogbn-arxiv, low-degree nodes exhibit higher calibration errors.
For issue (i), we note that besides the logits and features mentioned in the prior work (Wang et al., 2021; Huang et al., 2022), Figure 1 demonstrates that node degree also affects the calibration error, aligning with the message-passing mechanism in GNNs. In some cases, a low degree can lead to worse calibration performance due to insufficient information passing. Conversely, overconfidence may occur when excessive (potentially misleading) information floods into a single node (Wang et al., 2022). Regarding issue (ii), it remains unclear how to develop an appropriate algorithm that incorporates multiple factors for calibration within the graph structure. To address these issues, we propose a novel calibration framework named Graph Ensemble Temperature Scaling (GETS). Our approach leverages both input ensemble and model ensemble strategies within a Graph Mixture-of-Experts (MoE) architecture to address the limitation of existing methods. By incorporating multiple calibration experts—each specializing in different aspects of the influential factors affecting calibration—GETS can adaptively select and combine relevant information for each node.
@@
  • We introduce a novel framework GETS that combines input and model ensemble strategies within a Graph MoE architecture. This enables robust node-wise calibration by integrating diverse inputs like logits, features, and degree embeddings.
  • GETS significantly reduces ECE across 10 GNN benchmark datasets, outperforming state-of-the-art methods like CaGCN and GATS, and demonstrating effectiveness across diverse graph structures.
  • GETS employs the sparse gating mechanism to select the top $k$ experts per node. This enables scalable calibration on large graphs with computational efficiency.
@@ 2.1 Uncertainty Calibration
Uncertainty calibration adjusts a model’s predicted probabilities to better reflect the true likelihood of outcomes. Methods for uncertainty calibration are generally classified into in-training and post-hoc approaches.
In-training calibration adjusts uncertainty during model training. Bayesian methods, such as Bayesian Neural Networks (BNNs) and variational inference, adjust uncertainty by placing distributions over model parameters (Gal & Ghahramani, 2016; Sensoy et al., 2018; Detommaso et al., 2022). Frequentist approaches, including conformal prediction (Zargarbashi et al., 2023; Huang et al., 2024; Deshpande et al., 2024) and quantile regression (Chung et al., 2021), are also employed to ensure calibrated outputs.
Despite these efforts, model misspecification and inaccurate inference often lead to miscalibration (Kuleshov et al., 2018). To address this, post-hoc methods recalibrate models after training. These can be divided into nonparametric and parametric approaches. Nonparametric methods, such as histogram binning (Zadrozny & Elkan, 2001), isotonic regression (Zadrozny & Elkan, 2002), and Dirichlet calibration (Kull et al., 2019), adjust predictions flexibly based on observed data (Naeini et al., 2014). Parametric methods assume a specific functional form and learn parameters that directly modify predictions. Examples include temperature scaling (Guo et al., 2017), Platt scaling (Platt et al., 1999), beta calibration (Kull et al., 2017) and so on.
@@ 3.1 Uncertainty Calibration on GNN
This paper addresses the calibration of common semi-supervised node classification tasks. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ represent the graph, where $\mathcal{V}$ is the set of nodes, and $\mathcal{E}$ is the set of edges. The adjacency matrix is denoted by $\mathbf{A}\in\mathbb{R}^{N\times N}$, with $N=|\mathcal{V}|$ being the number of nodes. Let $\mathcal{X}$ be the input space and $\mathcal{Y}$ be the label space, with $y_{i}\in\mathcal{Y}=\{1,\dots,K\}$, where $K\geq 2$ is the number of classes. Let $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$ be the input feature and label. A subset of nodes with ground-truth labels $\mathcal{L}\subset\mathcal{V}$ is selected for training. The goal of semi-supervised node classification is to infer the labels of the unlabeled nodes, $\𝒰\𝒱\mathcal{U}=\mathcal{V}\backslash\mathcal{L}$. The node-wise input feature matrix and the label vector are denoted as $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{N}]^{\top}$ and $\mathbf{Y}$, respectively.
A GNN model $f_{\theta}$ is trained to address this problem by considering the node-wise features $\{\mathbf{x}_{i}\}_{i:v_{i}\in\mathcal{L}}$ and $\mathbf{A}$, where $\theta$ represents the learnable parameters. For a given node $i$, the logit output of the GNN is represented as $\mathbf{z}_{i}=f_{\theta}(\mathbf{x}_{i},\mathbf{A})=[z_{i,1},\dots,z_{i,K}]^{\top}$, where $K$ is the total number of classes. The predicted label for node $i$ is given by $\hat{y}_{i}=\arg\max_{k\in[1,K]}z_{i,k}$, selecting the class with the highest logit value. The confidence or predicted probability for node $i$ is then defined as $\hat{p}_{i}=\max_{k\in[1,K]}\operatorname{softmax}(z_{i,k})$, where $\operatorname{softmax}(z_{i,k})$ transforms the logits into probability scores for each class $k$. We define $f_{\theta}$ to be perfectly calibrated as (Guo et al., 2017; Chung et al., 2021):
@@
If $\text{CE}(f_{\theta})=0$, then $f_{\theta}$ is perfectly calibrated. However, direct computation of the calibration error from the model outputs is challenging in practice. Therefore, the expected calibration error (ECE) is proposed to approximate CE by discretization (binning) and summation, where the model outputs are partitioned into $B$ intervals (bins) with binning strategy =$\mathcal{B}=\{I_{1},I_{2},\dots,I_{B}\}$. A common binning strategy $\mathcal{B}^{0}$ divides bins by sorting confidence scores $\hat{p}_{i}$ in ascending order, ensuring each bin $I^{0}_{j}$ contains approximately the same number of nodes, i.e., $|I^{0}_{j}|\approx\frac{N}{B}$.
The ECE is defined as the weighted average of the absolute difference between the accuracy and confidence in each bin, denoted as $Acc(I_{j})$ and $Conf(I_{j})$:
@@ 3.2 Node-wise Parametric Calibration
TS is the classical parametric calibration approach, adjusts model confidence by scaling logits with a temperature $T>0$ before softmax, smoothing outputs without affecting accuracy. The temperature $T$ controls the smoothness of predicted probabilities: $T=1$ uses logits directly, while increasing $T$ spreads probabilities, improving calibration in overconfident models.
@@ 4.1 Ensemble Strategy Meets Node-wise Calibration
Ensemble strategy are widely recognized for improving model robustness and calibration by aggregating predictions from multiple models. For example, in the context of ensemble temperature scaling (ETS) (Zhang et al., 2020), each model learns its own temperature parameter $T_{m}$ to calibrate its predictions. The ensemble strategy combines the calibrated outputs from multiple calibration models, averaging their predictions to produce a more reliable and better-calibrated result. The ensemble weighting is formulated as a weighted sum over temperature-scaled logits (Zhang et al., 2020): $T(z)=\sum_{m=1}^{M}w_{m}\cdot t_{m}(z)$, where $w_{m}$ are non-negative coefficients summing to 1, and $t_{m}(z)$ represents the value of temperature of the $m$-th model TS model trained on logit $z$, with $M$ being the number of models.
@@
We propose to learn node-wise ensemble strategy for the GNN calibration task. To incorporate the different sources of information, we ensemble based on not only the calibration models but also diverse inputs, including logits $z_{i}$, node feature $\mathbf{x}_{i}$ and node degrees $d_{i}$, as well as their combination as the calibration model inputs. We use $g_{m}$ to represent the $m$-th GNN calibration model that outputs the calibrated logits. Due to the monotone-preserving property of TS, the weighted summation of the logits still keeps the invariant accuracy output. For each node $i$, the calibrated logits $z_{i}^{\prime}$ are computed as a weighted sum of the outputs from multiple calibration models:
@@
with $w_{i,m}\geq 0$ are the node-specific ensemble weights for node $i$ and model $m$, satisfying $\sum_{m=1}^{M}w_{i,m}=1$. We further point out that in our expert setting, the input of each expert only contains one element in $\mathbf{x}_{i},d_{i},z_{i}$. This gives reasonable learning tasks for different experts and rules out the problem arising in concatenating vectors of different input spaces. Each calibration model $g_{m}$ can focus on different aspects of the calibration process by utilizing various combinations of inputs. For example, one model may calibrate based solely on the logits, another may incorporate node features, and another may consider the node degrees. Specifically, $w_{i,m}$ can therefore be determined by a gating function $G(\cdot)$ that outputs a probability distribution over the calibration models for each node:
@@
where $\mathbf{W}_{g}$ and $\mathbf{W}_{n}$ represent the parameters of the gating function $G(\cdot)$, which will be introduced later on. We ensemble the feature inputs by concatenating different combinations of inputs into $g_{m}(\cdot)$. The key component remains unknown is how we should do the information aggregation and obtain the gating function to select models.
We include logits $z$, node features $\mathbf{x}$, and degree (embeddings) $d$ to capture key factors influencing calibration. Including logits is crucial because they contain the raw prediction information before applying softmax, reflecting the model’s confidence. Incorporating node features allows the model to address individual calibration needs, as feature-based methods binning methods have successfully captured unique sample point characteristics in calibration task (Huang et al., 2022). Adding degree embeddings tackles structural disparities since our experiments show that degree imbalance leads to varying calibration errors among nodes. By combining these inputs, our ensemble leverages multiple information sources, resulting in a more robust and effective calibration strategy.

### v2

@@ Abstract
Graph Neural Networks (GNNs) deliver strong classification results but often suffer from poor calibration performance, leading to overconfidence or underconfidence. This is particularly problematic in high-stakes applications where accurate uncertainty estimates are essential. Existing post-hoc methods, such as temperature scaling, fail to effectively utilize graph structures, while current GNN calibration methods often overlook the potential of leveraging diverse input information and model ensembles jointly. In this paper, we propose Graph Ensemble Temperature Scaling (GETS), a novel calibration framework that combines input and model ensemble strategies within a Graph Mixture-of-Experts (MoE) architecture. GETS integrates diverse inputs, including logits, node features, and degree embeddings, and adaptively selects the most relevant experts for each node’s calibration procedure. Our method outperforms state-of-the-art calibration techniques, reducing expected calibration error (ECE) by $\geq 25\%$ across 10 GNN benchmark datasets. Additionally, GETS is computationally efficient, scalable, and capable of selecting effective input combinations for improved calibration performance. The implementation is available at https://github.com/ZhuangDingyi/GETS/.
@@ 1 Introduction
Graph Neural Networks (GNNs) have emerged as powerful tools for learning representations in numerous real-world applications, including social networks, recommendation systems, biological networks, and traffic systems, achieving state-of-the-art performance in tasks like node classification, link prediction, and graph classification (Kipf & Welling, 2016; Veličković et al., 2017). Their ability to capture complex relationships and dependencies makes them invaluable for modeling and predicting interconnected systems (Fan et al., 2019; Wu et al., 2019; Scarselli et al., 2008; Wu et al., 2022; 2021; Jiang* et al., 2024; Luo et al., 2024).
However, ensuring the reliability and trustworthiness of GNN prediction remains a critical challenge, especially in high-stakes applications where decisions can have significant consequences (e.g. human safety). Consider GNN classification as an example, one key aspect of refining model reliability is uncertainty calibration, which aims to align the predicted probabilities with the true likelihood of outcomes (Guo et al., 2017). In short, well-calibrated models provide confidence estimates that reflect real-world probabilities, which is essential for risk assessment and informed decision-making (Zhuang et al., 2023).
Calibration methods are typically applied post-hoc and can be broadly classified into two categories: nonparametric and parametric. Nonparametric methods, such as histogram binning and isotonic regression, adjust confidence estimates based on observed data distributions without assuming a specific functional form (Zadrozny & Elkan, 2001; Naeini et al., 2015). While these methods allow flexibility in modeling, they often require large calibration datasets and can struggle with high-dimensional or unevenly distributed graph data, making them less practical for graphs with vast numbers of nodes. In contrast, parametric methods, including temperature scaling (TS) and Platt scaling, assume a functional relationship to adjust model outputs (Guo et al., 2017; Platt et al., 1999). Among these, TS has gained popularity due to its simplicity and effectiveness in calibrating models by smoothing predicted probabilities. This method is scalable and easy to implement, especially for large-scale graphs. Recent work has moved beyond the uniform adjustments applied to all data in classic TS, focusing on learning node-specific calibration parameters to capture individual characteristics that affect graph calibration performance (Tang et al., 2024; Hsu et al., 2022).
@@
Given these challenges, we aim to address the following issues: (i) What are the key factors influencing the calibration task for GNN classification? (ii) How can we effectively integrate these factors, which may exist across different spaces and scales, to perform GNN calibration?
Figure 1: Expected calibration error (ECE), see Equation 3, for a well-trained CaGCN model (Wang et al., 2021). ECE is measured by grouping nodes based on degree ranges rather than predicted confidence. Datasets are sorted by average degree $\frac{2|\mathcal{E}|}{|\mathcal{V}|}$ as a measure of connectivity (see Table 1), from low to high. In lower-connectivity datasets like Pubmed and Cora-full, high-degree nodes tend to show larger calibration errors, whereas in more connected datasets like Photo and Ogbn-arxiv, low-degree nodes exhibit higher calibration errors.
For issue (i), we observe that, in addition to the logits and features discussed in prior work (Wang et al., 2021; Huang et al., 2022), Figure 1 demonstrates that node degree influences the calibration error, consistent with the message-passing mechanism in GNNs. In some cases, a low node degree can result in worse calibration performance due to insufficient information propagation. Conversely, excessive (and potentially misleading) information may flood a single node, leading to overconfidence in its predictions (Wang et al., 2022). Regarding issue (ii), it remains unclear how to develop an effective algorithm that incorporates multiple factors for calibration within the graph structure. To address these challenges, we propose a novel calibration framework called Graph Ensemble Temperature Scaling (GETS). Our approach integrates both input ensemble and model ensemble strategies within a Graph Mixture-of-Experts (MoE) architecture, thereby overcoming the limitations of existing methods. By incorporating multiple calibration experts—each specializing in different aspects of the influential factors affecting calibration—GETS can adaptively select and combine relevant information for each node.
@@
  • We introduce a novel framework, GETS, which combines input and model ensemble strategies within a Graph MoE architecture. This enables robust node-wise calibration by integrating diverse inputs, such as logits, features, and degree embeddings.
  • GETS significantly reduces Expected Calibration Error (ECE) across 10 GNN benchmark datasets, outperforming state-of-the-art methods like CaGCN and GATS, and demonstrating effectiveness across diverse graph structures.
  • GETS employs a sparse gating mechanism to select the top $k$ experts for each node, enabling scalable calibration on large graphs while maintaining computational efficiency.
@@ 2.1 Uncertainty Calibration
Uncertainty calibration adjusts the predicted probabilities of the model to more accurately reflect the true likelihood of outcomes. Calibration methods are generally classified into in-training and post-hoc approaches.
In-training calibration adjusts uncertainty during model training. Bayesian methods, such as Bayesian Neural Networks (BNNs) and variational inference, achieve this by placing distributions over model parameters (Gal & Ghahramani, 2016; Sensoy et al., 2018; Detommaso et al., 2022). Frequentist approaches, including conformal prediction (Zargarbashi et al., 2023; Huang et al., 2024; Deshpande et al., 2024) and quantile regression (Chung et al., 2021), are also employed to ensure well-calibrated outputs.
Despite these efforts, model misspecification and inaccurate inference often lead to miscalibration (Kuleshov et al., 2018). To address this, post-hoc methods recalibrate models after training. These can be divided into nonparametric and parametric approaches. Nonparametric methods, such as histogram binning (Zadrozny & Elkan, 2001), isotonic regression (Zadrozny & Elkan, 2002), and Dirichlet calibration (Kull et al., 2019), adjust predictions flexibly based on observed data (Naeini et al., 2014). Parametric methods assume a specific functional form and learn parameters that directly modify predictions. Examples include temperature scaling (Guo et al., 2017), Platt scaling (Platt et al., 1999), beta calibration (Kull et al., 2017), and so on.
@@ 3.1 Uncertainty Calibration on GNN
This paper addresses the calibration of common semi-supervised node classification tasks. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ represent the graph, where $\mathcal{V}$ is the set of nodes and $\mathcal{E}$ is the set of edges. The adjacency matrix is denoted by $\mathbf{A}\in\mathbb{R}^{N\times N}$, where $N=|\mathcal{V}|$ represents the number of nodes. Let $\mathcal{X}$ be the input space and $\mathcal{Y}$ be the label space, with $y_{i}\in\mathcal{Y}=\{1,\dots,K\}$, where $K\geq 2$ is the number of classes. Let $\mathbf{X}\in\mathcal{X}$ and $\mathbf{Y}\in\mathcal{Y}$ denote the input features and labels, respectively. A subset of nodes with ground-truth labels, $\mathcal{L}\subset\mathcal{V}$, is selected for training. The goal of semi-supervised node classification is to infer the labels of the unlabeled nodes, $\𝒰\𝒱\mathcal{U}=\mathcal{V}\backslash\mathcal{L}$. The node-wise input feature matrix and the label vector are denoted by $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{N}]^{\top}$ and $\mathbf{Y}$, respectively.
A GNN model $f_{\theta}$ is trained to address this problem by considering the node-wise features $\{\mathbf{x}_{i}\}_{i:v_{i}\in\mathcal{L}}$ and the adjacency matrix $\mathbf{A}$, where $\theta$ represents the learnable parameters. For a given node $i$, the logit output of the GNN is represented as $\mathbf{z}_{i}=f_{\theta}(\mathbf{x}_{i},\mathbf{A})=[z_{i,1},\dots,z_{i,K}]^{\top}$, where $K$ is the total number of classes. The predicted label for node $i$ is given by $\hat{y}_{i}=\arg\max_{k\in[1,K]}z_{i,k}$, selecting the class with the highest logit value. The confidence or predicted probability for node $i$ is then defined as $\hat{p}_{i}=\max_{k\in[1,K]}\operatorname{softmax}(z_{i,k})$, where $\operatorname{softmax}(z_{i,k})$ transforms the logits into probability scores for each class $k$. We define $f_{\theta}$ to be perfectly calibrated as (Guo et al., 2017; Chung et al., 2021):
@@
If $\text{CE}(f_{\theta})=0$, then $f_{\theta}$ is perfectly calibrated. However, direct computation of the calibration error from the model outputs is challenging in practice. Therefore, the ECE is proposed to approximate CE by discretization (binning) and summation, where the model outputs are partitioned into $B$ intervals (bins) with binning strategy =$\mathcal{B}=\{I_{1},I_{2},\dots,I_{B}\}$. A common binning strategy $\mathcal{B}^{0}$ divides bins by sorting confidence scores $\hat{p}_{i}$ in ascending order, ensuring each bin $I^{0}_{j}$ contains approximately the same number of nodes, i.e., $|I^{0}_{j}|\approx\frac{N}{B}$.
The ECE is defined as the weighted average of the absolute difference between the accuracy and confidence in each bin, denoted by $Acc(I_{j})$ and $Conf(I_{j})$:
@@ 3.2 Node-wise Parametric Calibration
TS is the classical parametric calibration approach, which adjusts model confidence by scaling logits with a temperature $T>0$ before softmax, smoothing outputs without affecting accuracy. The temperature $T$ controls the smoothness of predicted probabilities: $T=1$ uses logits directly, while increasing $T$ spreads probabilities, improving calibration in overconfident models.
@@ 4.1 Ensemble Strategy Meets Node-wise Calibration
Ensemble strategies are widely recognized for improving model robustness and calibration by aggregating predictions from multiple models. For example, in the context of ensemble temperature scaling (ETS) (Zhang et al., 2020), each model learns its own temperature parameter $T_{m}$ to calibrate its predictions. The ensemble strategy combines the calibrated outputs from multiple calibration models, averaging their predictions to produce a more reliable and better-calibrated result. The ensemble weighting is formulated as a weighted sum over temperature-scaled logits (Zhang et al., 2020): $T(z)=\sum_{m=1}^{M}w_{m}\cdot t_{m}(z)$, where $w_{m}$ are non-negative coefficients summing to 1, and $t_{m}(z)$ represents the value of temperature of the $m$-th model TS model trained on logit $z$, with $M$ being the number of models.
@@
We propose to learn a node-wise ensemble strategy for the GNN calibration task. To incorporate the different sources of information, we ensemble based on not only the calibration models but also diverse inputs, including logits $z_{i}$, node feature $\mathbf{x}_{i}$ and node degrees $d_{i}$, as well as their combination as the calibration model inputs. We use $g_{m}$ to represent the $m$-th GNN calibration model that outputs the calibrated logits. Due to the monotone-preserving property of TS, the weighted summation of the logits still keeps the invariant accuracy output. For each node $i$, the calibrated logits $z_{i}^{\prime}$ are computed as a weighted sum of the outputs from multiple calibration models:
@@
where $w_{i,m}\geq 0$ are the node-specific ensemble weights for node $i$ and model $m$, satisfying $\sum_{m=1}^{M}w_{i,m}=1$. We further point out that in our expert setting, the input of each expert only contains one element in $\mathbf{x}_{i},d_{i},z_{i}$. This gives reasonable learning tasks for different experts and rules out the problem arising in concatenating vectors of different input spaces. Each calibration model $g_{m}$ can focus on different aspects of the calibration process by utilizing various combinations of inputs. For example, one model may calibrate based solely on the logits, another may incorporate node features, and another may consider the node degrees. Specifically, $w_{i,m}$ can therefore be determined by a gating function $G(\cdot)$ that outputs a probability distribution over the calibration models for each node:
@@
where $\mathbf{W}_{g}$ and $\mathbf{W}_{n}$ represent the parameters of the gating function $G(\cdot)$, which will be introduced later on. We ensemble the feature inputs by concatenating different combinations of inputs into $g_{m}(\cdot)$. The key component that remains unknown is how we should do the information aggregation and obtain the gating function to select models.
We include logits $z$, node features $\mathbf{x}$, and degree (embeddings) $d$ to capture key factors influencing calibration. Including logits is crucial because they contain the raw prediction information before applying softmax, reflecting the model’s confidence. Incorporating node features allows the model to address individual calibration needs, as feature-based methods binning methods have successfully captured unique sample point characteristics in calibration task (Huang et al., 2022). Adding degree embeddings tackles structural disparities since our experiments show that degree imbalance leads to varying calibration errors among nodes. Note that degrees are the foundation for graph structure information but we also explore other structural features as inputs in Appendix A.9. By combining these inputs, our ensemble leverages multiple information sources, resulting in a more robust and effective calibration strategy.